### v1

# Parameter Generation of Quantum Approximate Optimization Algorithm with Diffusion Models

Abstract
Quantum computing presents a compelling prospect for revolutionizing the field of combinatorial optimization, in virtue of the unique attributes of quantum mechanics such as superposition and entanglement. The Quantum Approximate Optimization Algorithm (QAOA), a variational hybrid quantum-classical algorithm, stands out as leading proposals to efficiently solve the Max-Cut problem, a representative example of combinatorial optimization. However, its promised advantages strongly rely on parameters initialization strategy, a critical aspect due to the non-convex and complex optimization landscapes characterized by low-quality local minima issue. Therefore, in this work, we formulate the problem of finding good initial parameters as a generative task in which the generative machine learning model, specifically denoising diffusion probalistic model (DDPM), is trained to generate high-performing initial parameters for QAOA. The diffusion model is capable of learning the distribution of high-performing parameters and then synthesizing new parameters closer to optimal ones. Experiments with various sized Max-Cut problem instances demonstrate that our diffusion process consistently enhance QAOA’s effectiveness compared to random parameters initialization. Moreover, our framework shows the possibility of training on small, classically simulatable problem instances, aiming at extrapolating to larger instances to reduce quantum computational resource overhead.
1 Introduction
Quantum computing is one of the major transformative technologies and presents an entirely new computational paradigm that has potential to achieve an exponential or polynomial advantage for classically intractable problems encompassing machine learning [1, 2, 3, 4], molecular dynamics [5, 6, 7, 8], and combinatorial optimization [9, 10, 11, 12, 13], etc. Although quantum computing is still in the noisy intermediate-scale quantum (NISQ) era [14] where quantum hardware is characterized by the limited qubit counts, inherent system noise, significant quantum gates error, and constrained qubit topology, the community have dedicated to exploring quantum algorithms tailored to NISQ machines.
Among these efforts, Quantum Approximate Optimization Algorithm (QAOA) [9, 10, 15, 16], revolving around the use of Parameterized Quantum Circuit (PQC) similarly to classical neural network, emerges as the foremost proposal and has exhibited their superiority in address complex and NP-Hard combinatorial optimization problems like the Max-Cut problem [9, 17, 18, 19]. QAOA combines a parameterized quantum state evolution that is performed on a NISQ device, with a classical optimizer that is used to find optimal parameters. Furthermore, QAOA advance spurs the exploration of practical applications such as circuit layout designs [20], wireless communication [21, 22], finance [23, 24] and so on.
Although QAOA shows immense practical significance in many fields, optimizing QAOA parameters poses a significant challenges [16, 25, 26] because the optimization objective is nonconvex with low-quality nondegenerate local minima of the cost landscape hindering the trainability of the algorithm. Parameters optimization focus on evolving the random initial parameters to the specific ones that can perform well on given tasks. Many approaches have been proposed for QAOA parameter optimization including gradient-based and derivative-free methods [27, 28, 29]. However, these methods usually require many measurement runs and consequently remain resource-expensive computations. Further, parameters initialization has been recognized as a crucial pipeline to promote the convergence to a potential solution within the parameter space, thereby facilitating more effective optimization without resorting to resource-expensive computations [30, 31]. Therefore, we impose that it is essential and inevitable to explore a high-quality parameter initialization strategy.
Diffusion model [32, 33, 34], an success generative machine learning method, has shown the significant advantage in image and visual generation. Recently, the denoising diffusion model has been utilized to generate quantum circuits for unitary compilation and entanglement generation tasks [35]. Taking a closer look at opimizing QAOA parameters and diffusion models, we are aware of the commonalities between diffusion-based image generation and QAOA parameters optimizaion process in the following aspects (illustrated in Figure 1 introduced by [36]): $\left(i\right)$ both the reverse process of diffusion model and QAOA parameters optimizaion can be regarded as transitions from random noise/initialization to specific distributions. $\left(ii\right)$ high-performing QAOA parameters and high-quality images can be degraded into simple distributions , like Gaussian distributions, by iteratively adding noise. To this end, motivated by the foregoing similarity, we delve into the synergy between the use of diffusion model and QAOA parameters initialization where the trained model can explore the evolution from random parameters to high-performing parameters. We posit that once the model trained, it can efficiently learn the distribution of high-performing parameters, and excel at generating a good initialization to speedup the optimization convergence. This innovative approach paves the way for harnessing the capabilities of both generative machine learning model and quantum computing to address some of the most formidable challenges in computational science.
The main contributions of our work are summarized as follows.
  • A denoising diffusion probalistic model is constructed and trained to generate high-performing initial parameters for QAOA. The proposed approach leverages the similarity between diffusion-based image and QAOA’s initial parameters generations, strengths of both quantum computing and generative machine learning model.
  • Extensive evaluations on various sized Max-Cut problem instances using the Xanadu Pennylane [37] quantum circuit simulator show that our proposed scheme notably outperforms random initialization strategy. Once the model is trained, the parameters generated can be leveraged to efficiently find high-quality initial parameters for unseen test instances whose sizes even are beyond the training dataset regime.
Figure 1: The left: illustrates the standard diffusion process in image generation. The right: denotes the parameter distribution of QAOA.
2 Preliminaries
2.1 Quantum Computing
[…]
The fundamental unit in quantum computation is the quantum bit (qubit). A qubit is a quantum state represented in a two-dimensional space and can be expressed as $\left|\psi\right\rangle=a\left|0\right\rangle+b\left|1\right\rangle$. Further, the quantum state represented by two qubits is constructed using the tensor product, i.e., $\left|a\right\rangle\otimes\left|b\right\rangle$ as a 4-dimensional vector. The quantum operations on qubits are quantum gates, as unitary matrices, transforming a quantum state to another and preserving the norm of the quantum state. The basic quantum gates can be split into two groups, single-qubit gates and two-qubit gates. The commonly used single-qubit gates are Pauli gates ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$, and $I$) and corresponding rotation gates, and Hadamard gate ($H$), which can respectively be denoted as the following unitary matrices,
$\sigma_{x}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ $\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ $\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (1)
$R_{x}\left(\theta\right)=\begin{pmatrix}\cos\frac{\theta}{2}&-i\sin\frac{\theta}{2}\\ -i\sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix}$ $R_{y}\left(\theta\right)=\begin{pmatrix}\cos\frac{\theta}{2}&-\sin\frac{\theta}{2}\\ \sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix}$ $R_{z}\left(\theta\right)=\begin{pmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{pmatrix}$ (2)
The notable two-qubit gate is the CNOT gate as follows
$CNOT=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$ (3)
2.2 Quantum Approximate Qptimization Algorithm for Max-Cut
The Quantum Approximate Optimization Algorithm (QAOA) is a variational quantum algorithm paradigm developed to address combinatorial optimization tasks on NISQ devices. It utilizes the problem-dependent parameterized quantum circuit involving alternate Hamiltonian and Mixer layers to prepare and evolve quantum states that encode potential solutions to the optimization problem at hand. By adjusting variational parameters and optimizing them using classical optimization techniques, QAOA aims to find near-optimal solutions efficiently, as illustrated in Figure 2. QAOA is particularly suited for Max-Cut problem and leverages the expressive power of quantum computation to explore large solution spaces more effectively than classical counterparts. Its efficacy lies in classical initialization and optimization of the parameters, balancing the quantum circuit depth, parameter space exploration to achieve good approximation solutions.
Figure 2: The specific workflow of QAOA.
In QAOA, the binary assignment combinatorial optimization is first encoded in a cost Hamiltonian $H_{c}$ by the mapping between classical binary decision variables $s_{i}\in\left\{0,1\right\},i=1,2,\dots,n$ and the eigenvalues of the quantum Pauli $\sigma_{z}$, where the ground energy eigenstate of $H_{c}$ corresponds to the solution of the combinatorial optimization problem. Second, the transverse field mixer Hamiltonian is constructed as $H_{m}={\textstyle\sum_{i=1}^{n}}\sigma_{x}^{i}$. Then, the uniform superposition initial state $\left|+\right\rangle^{\otimes n}$ is prepared by performing Hadamard gates on all qubits with an all-zeroes state. Next, a variational quantum state $\left|\psi\left(\gamma_{1},\dots,\gamma_{p},\beta_{1},\dots,\beta_{p}\right)\right\rangle$ is prepared by employing alternate Hamiltonian and Mixer layers $e^{-i\gamma_{i}H_{c}}$ and $e^{-i\beta_{i}H_{m}}$, $i=1,2,\dots,p$ as follows
$\left|\psi\left(\gamma_{1},\dots,\gamma_{p},\beta_{1},\dots,\beta_{p}\right)\right\rangle=e^{-i\beta_{p}H_{m}}e^{-i\gamma_{p}H_{c}}\cdots e^{-i\beta_{1}H_{m}}e^{-i\gamma_{1}H_{c}}\left|+\right\rangle^{\otimes n}$ (4)
Finally, the selected classical optimizer is applied to vary parameters $\gamma_{i}$ and $\beta_{i}$, $i=1,2,\dots,p$ to minimize the cost function
$C\left(\gamma_{1},\dots,\gamma_{p},\beta_{1},,\dots,\beta_{p}\right)=\left\langle\psi\left(\gamma_{1},\dots,\gamma_{p},\beta_{1},\dots,\beta_{p}\right)\right|H_{c}\left|\psi\left(\gamma_{1},\dots,\gamma_{p},\beta_{1},\dots,\beta_{p}\right)\right\rangle$ (5)
For the Max-Cut problem, given a graph $G=\left(V,E\right)$ where $V$ is the set of nodes and $E$ is the set of edges, the goal of Max-Cut is to partition the set of nodes $V$ into two disjoint subsets such that the total weight of edges connecting the two subsets is maximized. The mathematical expression of the Max-Cut is formulated as follows
$\underset{\vec{s}}{\min}\sum_{i,j\in Vand\left(i,j\right)\in E}-w_{ij}s_{i}s_{j}$ (6)
where $s_{k}\in\left\{-1,1\right\}$ and the weight $w_{ij}$ of edges is set to $1$ for the unweighted Max-Cut problem. To apply QAOA to the Max-Cut, the above objective can be encoded in the following problem Hamiltonian by mapping binary variables $s_{k}$ onto the eigenvalues of the Pauli $\sigma_{z}$
$\min\sum_{i,j\in Vand\left(i,j\right)\in E}-w_{ij}\sigma_{z}^{i}\sigma_{z}^{j}$ (7)
Thus, minimizing the objective of the Max-Cut is equivalently transformed into obtaining the ground state energy of the problem Hamiltonian.
2.3 Denoising Diffusion Probalistic Model
Diffusion model is a class of generative models traced back to non-equilibrium thermodynamics [38, 39]. The model aims at progressively removing noise from inputs and generate clear images. The representative work, denoising diffusion probalistic model (DDPM) [32], refine diffusion model with a training paradigm characterized by forward and reverse processes in a multi-step chain indexed by timesteps. The forward process, dubbed diffusion process, is characterized as progressive adding noise as well as the denoising process is regarded as a gradual eliminating noise process.
In diffusion process, given an original sample $x_{0}$, Gaussian noise is progressively added for $T$ steps and further the noisy samples $x_{1},x_{2},\dots,x_{T}$ are obtained. Based on the reparameterization rule, the mathematical expression of this process can be written as follows,
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon\quad or\quad q\left(x_{t}|x_{0}\right)=\mathcal{N}\left(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},\left(1-\bar{\alpha}_{t}\right)\epsilon^{2}\right)$ (8)
where $\bar{\alpha}_{1},\bar{\alpha}_{2},\dots,\bar{\alpha}_{T}$ are a series of constants decreasing with $t$ increasing.
In denoising process, beginning with the timestep $T$, taking a sample $x_{t}$ and the current timestep $t$ as inputs, the noise prediction network in the model can evaluate the noise and eliminate it from $x_{t}$ to generate a new sample $x_{t-1}$. Denoising process repeats the process until the original sample is restored. Mathematically, the denoising process can be formulated as follows
$\mathbf{x}_{t-1}\sim\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right)$ (9)
where $\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)$ is the noise prediction network and $\bar{\alpha}_{t}=\alpha_{t}\dots\alpha_{1}$.
3 Methodologoy
Due to the lack of publicly available data sets, we first construct the dateset consisting of various Max-Cut problems and the approximate optimal parameters of QAOA. Then, in the later part, we detailed the QAOA initial parameters diffusion model preparation.
3.1 Dataset Generation
We first construct synthetic regular graphs comprising 3500 instances with node sizes $N$ ranging from $4$ to $8$, and the probability $p$ of having an edge between two nodes varying between $0.3$ and $0.75$, illustrated in Figure 3. Next, each of the above graphs is inputted into QAOA for the Max-cut problem. Further, to maintain the consistency of the dimension of the diffusion model input and output, we fixed the layers of QAOA to $3$, which contains $6$ parameters $\mathbf{\vec{\gamma}}=\left(\gamma_{1},\gamma_{2},\gamma_{3}\right)$ and $\mathbf{\vec{\beta}}=\left(\beta_{1},\beta_{2},\beta_{3}\right)$. The QAOA starts with random initial parameters of $\mathbf{\vec{\gamma}}$ and $\mathbf{\vec{\beta}}$, and performs optimization process $500$ iterations. However, since the inherently complex optimization landscape of the QAOA algorithm [16, 25], the final parameters optimized may have a relatively large gap compared to optimal ones. Therefore, we devise the multi-start strategy to find alternative parameters closer to optimal values in which parallel optimization processes are undergoed with multiple random initial parameters. The parameters corresponding to the lowest cost function are recorded as the input of the proposed diffusion model.
Figure 3: The left: illustrates the graphs count distribution with nodes size varying. The right: denotes the graphs count distribution with various edge probabilities.
3.2 Model Preparation
In the pursuit of enhancing the application of QAOA to the max-cut problem, the proposed approach involves the integration of the DDPM model and QAOA parameters initialization. We first reshape the training data into the tensor representation. Next, as illustrated in Figure 4, the model is built with two process, named parameters diffusion and generation. In parameters diffusion process, to train the model, we corrupt the samples in the training set. Then, the model learns to predict the noise of the present step, aiming at subtracting it from the input sample. The noise prediction model is constructed including the encoding module and the step-embedding module. As illustrated in Figure 5, the encoding module involving $4$ full-connected layers with ReLU nonlinearities to encode the input parameters as well as the step-embedding module is equiped with $3$ embedding layers to iterative transform the time-step into the latent representation.
Figure 4: The top: illustrates the parameters diffusion. The bottom: denotes the parameters generation.
Figure 5: The architectures of the Encoder and Step-Embedding Modules.
Specifically, adapting from the work [32], the training procedure of the noise prediction network $\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)$ is characterized in Algorithm 1
Algorithm 1 The Noise prediction Model Training
1:  repeat
2:  $\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)$
3:  $t\sim\operatorname{Uniform}(\{1,\ldots,T\})$
4:  $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
5:  Take gradient descent step on $\nabla_{\theta}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},t\right)\right\|^{2}$
6:  until converged
where $\mathcal{N}(\mathbf{0},\mathbf{I})$ represents Gaussian noise distribution, $\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ is the noisy sample in $t$ time-step and $\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},t\right)$ predicts the present noise. And the parameters generation/sampling is characterized in Algorithm 2
Algorithm 2 Parameters Generation or Sampling
1:  $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:   for $t=T,\ldots,1\text{ do }$
3:    $t\sim\operatorname{Uniform}(\{1,\ldots,T\})$
4:    $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$ , else $\mathbf{z}=\mathbf{0}$
5:    $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right)+\sigma_{t}\mathbf{z}$
6:  end for
7:  return $x_{0}$
4 Experiment
In this section, we conduct experiments to validate the effectiveness of diffuion-based QAOA initial parameters generation.
4.1 Experiment Setup
Implementation Detail. For the noise prediction model, we set the input dimension of the encoder module to $6$ and the input dimension of the step-embedding module to $100$. The model is trained 100 epochs using a batch size of $50$ before examining it on test set. Furthermore, we use the Adam optimizer to optimize the model with the learning rate $lr=10^{-3}$, and no weight decay or learning rate scheduling is used.
Dataset and Baseline Model. We first randomly generate $50$ test graphs for the Max-Cut problem with different node sizes and edge probability within the training dataset regime to calculate the improvement of the attainable lowest energy/cost achieved by the diffusion-based parameters initialization over the baseline, namely, random initialization. We future promise to iteratively optimize parameters 100 steps after the above initialization strategies. The results are presented in Figure. 6, where the orange line is the attainable lowest energy for the proposed initialization, and the blue line for random initialization. The results show that the proposed initialization most outperforms the baseline and attains lower energies. The average improvements for different node sizes compared to the baseline are listed in Table 1. For clarity, we randomly generate 5 graphs with various node sizes to show the convergence of the proposed work against the baseline, as illustrated in Figure 7. Moreover, we randomly generate 8 graphs varying nodes size from 9 to 16 for the Max-Cut problem and demonstrate that the proposed initialization strategy can be effectively extrapolated to larger instances and enhance QAOA’s effectiveness compared to the baseline, as illustrated in Figure 8
Figure 6: Comparison of the lowest energy between the diffusion-based parameters initialization and the baseline.
Table 1: The average improvements of our proposed method against the baseline
Node size vs random initialization
4 1.16x
5 1.09x
6 1.13x
7 1.04x
8 1.14x
Figure 7: Cost convergence of the proposed method and the baseline.
Figure 8: Comparison of the diffusion-based parameters initialization and the baseline for larger instances.
5 Conclusion
In conclusion, delving into the commonalities between the diffusion-based generation process and QAOA’s parameters training, we present a significant advancement in combining the Denoising Diffusion Probalistic Model (DDPM) with the Quantum Approximate Optimization Algorithm (QAOA), specifically targeting the Max-Cut problem. To show the potential of the DDPM in enhancing the initialization process of QAOA parameters, we conduct extensive experiments. The results demonstrates that the integration of the DDPM with QAOA can improve the attainable lowest energy of QAOA compared to the baseline, Moreover, we also show that the proposed scheme has a capacity of generalizing the larger Max-Cut problem instances beyond the training data regime, consequently reduce quantum computational resource overhead. This research establishes the foundation for future explorations in quantum computing, highlighting the diffusion model as a crucial instrument in this rapidly evolving field.
6 Future Works
The current work offers multiple extensions. On the one hand, we may generalize QAOA parameters initialization to the partial or entire parameters initialization in the quantum neural networks. Moreover, we expect that the parameters sampling from the distribution learning by the diffusion model may mitigate the barren plateaus problem in the PQCs training. On the other hand, perhaps the more complex diffusion model such as U-Net’s cross-attention maps can future improve our method.
7 Acknowledgement
This work is supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent No.2022ZB139.